# Atomistic Framework for Glassy Polymer Viscoelasticity Across 20 Frequency Decades

### v1

@@ Abstract
Glassy polymers are central to engineering applications, yet their viscoelastic response over broad frequency and temperature ranges remains difficult to characterize. We extend non-affine deformation theory by incorporating a time-dependent memory kernel within the Generalized Langevin Equation for atomistic non-affine motions, yielding frequency-dependent mechanical response. Applied to poly(methyl methacrylate) (PMMA), the method captures the shear modulus and relaxation spectrum across more than twenty decades in frequency, from hundreds of terahertz down to the millihertz regime, thus bridging polymer mechanics from ordinary to extreme scales. Our predictions show quantitative consistency with independent estimates from oscillatory-shear molecular dynamics, Brillouin scattering, ultrasonic spectroscopy, Split-Hopkinson testing, and dynamic mechanical analysis (DMA), demonstrating a unified theoretical–computational route for multiscale characterization of polymer glasses.
Glassy polymers are widely used in structural and protective components, where their mechanical response governs performance and durability colmenero_2015polymers; wang_2024glassy, yet their viscoelastic behavior over a broad frequency range, covering molecular processes as well as engineering applications, remains poorly characterized. This shortcoming reflects the challenge of describing glassy polymer viscoelasticity consistently across the wide range of relevant timescales, both theoretically and experimentally. At high frequency level (GHz to THz), molecular simulations of glassy polymers have yielded detailed insight into relaxation dynamics near the glass transition, but at low frequency regimes ($\sim$ Hz) relevant to static load-bearing applications astm_2013standard; lakes2009viscoelastic; lakes_2017viscoelastic, viscoelasticity is often described by phenomenological models, in which the polymer’s broad spectrum of relaxation processes is compressed into adjustable parameter sets that represent effective modes rather than distinct molecular mechanisms tschoegl2012phenomenological. As a consequence, establishing a direct molecular-to-macroscopic link in the mechanical response of polymeric glasses remains challenging.
There are different experimental techniques available to probe the mechanical response of glassy polymers in various accessible frequency ranges. In the low-frequency regime (mHz-Hz), dynamic mechanical analysis (DMA) menard2020dynamic is a popularly used approach, where a sinusoidal deformation is applied and the corresponding response is measured. The internal friction of the material causes a phase lag between stress and strain, from which the storage modulus G’ and loss modulus G” are obtained Zaccone_book. DMA is an established technique to probe application-relevant conditions, construct long-term material behavior via time–temperature superposition, and assess how factors such as strain rate influence relaxation processes Wyss_PhysRevLett.98.238303; Sun_PhysRevB.111.184107; MULLIKEN20061331. As the probed frequency increases, available experimental methods transition from contact-based to non-contact and non-destructive. Dynamic contact-based mechanical testing, such as Split-Hopkinson pressure bars ($10^{3}$ - $10^{4}$ Hz) or ballistic testing (about $10^{6}$ Hz), determine material properties under high strain-rates, comparable to impacts and explosions. From MHz frequencies upward, only non-contact techniques are available, for example, ultrasonic methods, where sound waves are sent through the material, to determine their elastic properties afifi2003ultrasonic. In the GHz regime, Brillouin light scattering spectroscopy has been widely exploited Merklein_2022. As an inelastic light-scattering process, Brillouin scattering originates from interactions with high-frequency acoustic phonons, whose dynamics govern the magnitude of the so-called frequency shift that depends directly on the elastic properties of the medium. Brillouin spectroscopy provides access to microscopic mechanical response Merklein_2022; prevot_2025_softmatter, which has also recently been proven to agree well with atomistic modeling approaches for polymeric materials delasoudas2025elasticity; lakes2009viscoelastic.
Numerical investigations have long addressed the determination of elastic moduli in disordered solids Barrat_1988; Leonforte_PhysRevB.72.224206; Procaccia_PhysRevE.93.063003; kriuchevskyi_2017numerical; Krief_PhysRevE.103.063307, with many employing the stress-fluctuation formalism to evaluate the static shear modulus wittmer_2013shear; Wittmer_PhysRevE.91.022107; Kriuchevskyi_PhysRevLett.119.147802. Most of these studies rely on coarse-grained models, where the shear modulus is computed from equilibrium stress fluctuations. In such systems, stress correlations decay extremely slow as the glass transition approaches, reflecting the growing structural relaxation time. However, molecular dynamics (MD) simulations involving oscillatory shear remain limited to only a few orders of frequency range (above THz), due to computational constraints, thus making it difficult to access experimentally relevant timescales. Several advanced methods have been proposed to bridge this gap, including activation–relaxation theory Dyre_PhysRevB.53.2171; Olsen_PhysRevLett.81.1031, parallel replica dynamics perez_2015parallel, and non-affine deformation theory Leonforte_silica; Zaccone_book; Scossa; vaibhav2024time; singh2025viscosity. In spite of this, current computational approaches are only able to describe the dynamic mechanical response only at extremely short time scales mukherji.
A promising microscopic framework to address these limitations is the non-affine lattice dynamics (NALD) approach, which connects atomic-scale motion to macroscopic elastic response elder; vaibhav2024time. In NALD, the motion of each atom is described through its non-affine displacement relative to the affine deformation imposed by external strain. These non-affine displacements arise from structural disorder—due to thermal fluctuations or local topological asymmetries—that break local centrosymmetry. Under deformation, each atom experiences a net unbalanced force, the affine force field, which drives additional non-affine motion Zaccone_book. The internal work associated with these relaxations contributes negatively to the total deformation free energy, thereby reducing the macroscopic shear modulus Scossa. This framework provides a microscopic basis for understanding the elastic response of amorphous materials over a wide frequency range, thus establishing a theoretical route to link the atomistic dynamics with the experimental mechanical response elder; vaibhav2024time.
In this work, we develop and apply the microscopic non-affine computation scheme that exploits the atomic-scale description of particle displacements under deformation to overcome the intrinsic timescale limitations of conventional methods. This framework enables the computation of mechanical responses at experimentally accessible frequencies, thereby bridging the mechanical response from DMA to Brillouin scattering to atomistic MD. We study poly(methyl methacrylate) (PMMA) below its glass transition temperature, that has been extensively studied as a model system for exploring the viscoelastic response and relaxation spectra both experimentally and through molecular simulations MULLIKEN20061331; richeton_2007_modeling; ionita_2015viscoelastic; hu_2016_experimental; Lu_1997_polymer; William_PhysRevApplied.18.064078. We use atomistic MD simulations to explore the characteristics of the underlying potential energy surface. Using the normal mode analysis, we obtain the vibrational density of states (vDOS) and the affine force field correlator, which, together, enable the evaluation of the complex shear modulus. A time-dependent memory kernel is introduced to capture the primary and secondary relaxation decays of the shear modulus in the frequency domain. Furthermore, the theoretical estimates are validated with DMA measurements consisting of temperature sweeps at four different frequency levels (0.1, 1, 10, and 100 Hz), as well as with other experimental data in the Brillouin regime from the literature. Overall, a unique and advantageous aspect of our study is a description of polymer viscoelasticity across timescales of conventional mechanical testing (DMA), high rate tests (Split-Hopkinson testing, ultrasonics, and Brillouin spectroscopy), and molecular dynamics, ultimately spanning twenty frequency decades across molecular processes to application-relevant timescales.
@@
For the numerical study, we have considered 64 polymeric chains (each containing 10 monomers) of PMMA in a periodic simulation box, making a system of $N=9920$ atoms. Check section S2 of Supplementary Information (SI) supp for further details. The atomic interactions are modeled using the General AMBER Force Field (GAFF), where bond, angle, and dihedral terms are represented by harmonic potentials. Electrostatic interactions are treated using the particle-particle particle-mesh (PPPM) method, with a real-space cut-off of $9$ Å. Lennard-Jones interactions are truncated at $9$ Å. Large-scale molecular dynamics simulations are carried out using LAMMPS LAMMPS to generate polymer configurations under fixed temperature and pressure conditions. The system temperature and pressure are controlled through the Nosé–Hoover thermostat and barostat, respectively, with the pressure maintained at zero throughout the simulation. The equations of motion are solved under periodic boundary conditions using the velocity-Verlet integrator with a timestep of 1 fs.
@@
In NALD, potential energy surface (PES) wales_2006potential is required to measure the mechanical response of atomistic systems. PES describes the total potential energy of the system as a function of the atomic positions, effectively defining the energy landscape in which the atoms reside. The first derivative of the PES with respect to atomic positions gives the forces acting on each atom, which govern the atomic motion and local stability of the system, including internal stresses. The second derivative, known as the Hessian or dynamical matrix, provides information about the curvature of the PES, which determines the vibrational modes and eigenfrequencies of the system. These vibrational properties are crucial for calculating the affine (infinite-frequency) elastic modulus as well as the non-affine corrections arising from local, disorder-induced atomic rearrangements. We diagonalize the Hessian matrix to obtain the eigenfrequencies and corresponding eigenmodes of the system that are used to calculate the normalized vibrational density of states $g(\omega)=\frac{1}{3N-3}\sum_{i}\delta(\omega-\omega_{i})$ and the affine force field correlator $\Gamma_{\alpha\beta\kappa\chi}(\omega)=\langle\hat{\Xi}_{\alpha\beta}\hat{\Xi}_{\kappa\chi}\rangle_{\omega\in[\omega_{k},\omega_{k}+\delta\omega_{k}]}$ (see section S1 of SI supp for definition), where $\hat{\Xi}_{\alpha\beta}(\omega)$ is the component of affine force field. The frequency-dependent elastic constants $C_{\alpha\beta\kappa\chi}(\Omega)$ can be obtained in terms of these quantities via the relation Zaccone_book; Scossa
@@
where $\Omega$ is the applied strain frequency, $C^{\mathrm{A}}_{\alpha\beta\kappa\chi}$ is the affine contribution to elasticity, $\rho$ is number density, and $\nu$ is the friction memory kernel. The integrand term accounts for non-affine effects that act as a negative correction to the quasi-static limit, arising from local asymmetries in the atomic potential energy landscape during macroscopically affine deformations, as typically observed in glasses as well as in noncentrosymmetric crystals cui.
The various components of the elastic constants can be determined using Eq. (1) depend on different components of the affine force field with the corresponding component of the affine elastic constant. Here, we focus on the shear modulus $G$, for which Eq. (1) can be rewritten in the form
@@
Here, $G^{*}(\Omega)$ represents the complex shear modulus and $G_{A}$ is the affine modulus. We separate the real and imaginary components of $G^{*}(\Omega)$ (SI supp, section S1) that correspond to the storage and loss moduli, respectively, which quantify the elastic and viscous responses of the system. Also, we absorb the mass term in the affine force correlator and friction memory kernel kriuchevskyi2020scaling; vaibhav2024time to get corresponding mass-scaled quantity.
For simplicity, a fixed value of memory kernel $\nu$} is generally used in Eq. (2), which is the only parameter adjusted to perfectly match the MD results for the infinite-frequency plateau, as done also in vaibhav2024time; elder and discussed in Section S1 of SI supp. In a Markovian process, friction is treated as a constant, as is typically assumed in molecular simulations using e.g. the Langevin thermostat. However, in real glassy materials, friction is more accurately described within a non-Markovian framework, where the friction kernel depends on time or the history of deformation. To capture such experimental conditions, we allow the friction parameter $\nu$ to vary and adopt a simple power-law decay milster_2024tracer; li2017computing in the time domain, ${\nu}(t)\sim{\nu}_{0}t^{\delta-1}$ with $\delta<1$. For our needs in the frequency domain, the corresponding Fourier transform gives ${\nu}(\Omega)={\nu}_{0}\Omega^{-\delta}$, indicating that the effective friction decreases at high frequencies and recovers the constant Markovian limit as $\tau\to 0$. We use this friction kernel in the NALD equations to calculate the shear modulus in order to compare with the experimental results.
@@
We employ the NALD method to calculate the viscoelastic response and compare the results with experimental measurements. For the NALD calculations, we require microscopic information, namely the vibrational density of states $g(\omega)$, the affine force correlation function $\Gamma(\omega)$, and the friction kernel $\nu$. In Fig. 1(a) and (b), we plot the vibrational density of states (vDOS) and the affine force field correlator $\Gamma(\omega)$, respectively, as functions of the normal mode frequency $(\omega)$ at room temperature, T=300T=300K. Since our system is at a finite temperature, the vDOS contains a significant fraction of imaginary modes (instantaneous normal modes, INMs) (conventionally represented on the negative branch of $\omega$) in addition to real modes keyes1994unstable; Laird; stratt1995instantaneous; Jack; vaibhav2025experimental. The imaginary INMs originate from the presence of saddle points in the potential energy landscape, whereas the real modes correspond to stable vibrational motions (e.g. phonons).
@@
We observe several peaks across the entire frequency spectrum. These features may arise from the characteristic bonding nature of PMMA, where certain bonding interactions contribute vibrational modes in the high-frequency range of the vDOS. Near zero frequency, on both the positive and negative branches, the vDOS increases linearly, as often observed in glasses and supercooled liquids Petry; Laird; palyulin2018parameter. This low-frequency regime shows only little variation with temperature (as shown in Fig. S1 of SI supp).
@@
In Fig. 1(b), the affine-force field correlator $\Gamma(\omega)$ is shown for the same temperature. The overall profile of $\Gamma(\omega)$ exhibits several peaks, similar to the vDOS, with peak intensities increasing modestly as the temperature decreases (SI supp, section S1). At high frequencies, around $550\,\mathrm{THz}$, a pronounced peak is observed; however, its influence on the elastic modulus is minimal. In contrast, at low frequencies, $\Gamma(\omega)$ follows the expected analytical scaling  Scossa; Milkus_2017 $\Gamma(\omega)\sim\omega^{2}$, which plays a dominant role in the determination of the modulus. This low-frequency behavior reflects the long-wavelength vibrational modes that couple strongly with the macroscopic deformation response.
@@
Using the information from the vibrational density of states $g(\omega)$ and the affine force field correlator $\Gamma(\omega)$, we calculate the viscoelastic modulus as a function of the external frequency $\Omega$ at different temperatures below the glass transition. An important additional quantity in this calculation is the friction memory kernel, which enters the equations in the power-law form milster_2024tracer; li2017computing $\nu(\Omega)=\nu_{0}\Omega^{-\delta}$ with parameters $\nu_{0}=5.2\times 10^{18}\,\mathrm{kg\,s^{-1}}$ and $\delta=0.35$. In the time domain, this form corresponds to the memory kernel given by $\nu(t)\sim\nu_{0}t^{(\delta-1)}$. We plot and discuss $\nu$ in the SI supp as Figure S4 shows $\nu(\Omega)$ as a function of the external frequency.
@@
In the determination of the modulus $G$, we perform the integration over both the real and imaginary branches of $\omega$, as shown in Eqs. (S.4, S.5) of SI supp. Since the system is relatively small, finite-size effects arise, leading to the presence of non-physical low-frequency modes. To address this, a fraction of the small-frequency modes below a cut-off frequency $\omega_{\rm min}$ is discarded during the integration, as the system cannot support propagating modes with frequencies lower than the speed of sound in the medium. The minimum frequency can be estimated as $\omega_{\rm min}=\frac{2\pi}{L}\sqrt{\frac{G}{\rho}}$, where $L$ is the size of the simulation box, $\rho$ is the mass density, and $G$ is the shear modulus vaibhav2024time. Hence, when performing the finite analytical integration for the modulus calculations, we incorporate all modes with $|\omega|<\omega_{\min}$ to account for the low-frequency contributions below the cut-off.
@@
We observed a secondary $\beta$ relaxation Ngai in the shear modulus in the frequency domain around 1 Hz. This is reflected in the time domain, where $G(t)$ decays at $\tau\sim 1$ s. Capturing such small frequencies in simulations is very challenging due to system size limitations and the computational cost of Hessian diagonalization. To overcome this, we use an analytical-theory extension in the very low-frequency regime, exploiting the known behavior of $\Gamma(\omega)$ and $g(\omega)$ Zaccone_book. Specifically, we fit $\Gamma(\omega)=a_{0}\,\omega^{2},\quad g(\omega)=a_{1}\,\omega,$ with $a_{0}=3.1\times 10^{-18}$ and $a_{1}=4.5\times 10^{-16}$ Figs. S2, S3 corresponding plot in the SI supp. We then evaluate the additional analytical contribution, obtaining the exact integral solution at finite integration limits for both the real and imaginary parts of the eigenfrequencies. In this regime, the friction kernel is found to be on the order of $\nu=5\times 10^{22}\,\mathrm{kg/s}$ across the entire deformation frequency range.
@@
Figure 2: Temperature dependence of shear storage modulus $G^{\prime}$ for PMMA calculated using NALD (close triangle) is compared with DMA experimental data (closed diamonds, at external frequency $\Omega=1$ Hz). The solid line represents the theoretical fit based on the equation described in SI supp, with a glass transition temperature of $T_{g}\approx 390~\mathrm{K}$. We also plot the affine modulus $G_{A}$ (open squares) as a function of temperature.
@@
In Fig. 2, we show the shear modulus $G^{\prime}$ serving as an order parameter that distinguishes the glassy and liquid phases. Here, $G^{\prime}$ is plotted as a function of temperature at a very low deformation frequency of 1 Hz. We compare the experimental DMA results (closed diamonds) with atomistic NALD calculations (close triangles), which show a reasonable agreement. We also fit the experimental data using the theoretical Zaccone-Terentjev expression for $G^{\prime}$ vs $T$ (details in SI supp) ZET, which captures the sharp decay near the glass transition temperature $T_{g}=390~\mathrm{K}$. We observed a slight decrease in the affine modulus $G_{A}$ near and above $T_{g}$ due to enhanced molecular mobility and reduced topological constraints.
@@
Figure 3: Low-frequency shear storage modulus $G^{\prime}$ as a function of external deformation frequency $\Omega$ at temperature $T=300$ K. The Solid line represents NALD calculations and symbols represent MD and experimental results that include DMA study in this work and other experiments from previous studies weishaupt1995pressure; Mikio_1995; afifi2003ultrasonic; afifi2003annealing; MOTT2008572; Sahraoui_1994; davies1963dynamic; moy_paul_2003; MULLIKEN20061331; christofl2021comprehensive.

### v2

# Atomistic Framework for Glassy Polymer Viscoelasticity Across Twenty Frequency Decades

@@ Abstract
Glassy polymers are central to engineering applications, yet their viscoelastic response over broad frequency and temperature ranges remains difficult to characterize. We extend non-affine deformation theory by incorporating a time-dependent memory kernel within the Generalized Langevin Equation for atomistic non-affine motions, yielding frequency-dependent mechanical response. Applied to poly(methyl methacrylate) (PMMA), the method captures the shear modulus and relaxation spectrum across more than twenty decades in frequency, from hundreds of terahertz to the millihertz regime, thus bridging polymer mechanics from ordinary to extreme scales. Our predictions show quantitative consistency with independent estimates from oscillatory-shear molecular dynamics, Brillouin scattering, ultrasonic spectroscopy, Split-Hopkinson testing, and dynamic mechanical analysis (DMA), demonstrating a unified theoretical–computational route for multiscale characterization of polymer glasses.
Glassy polymers are essential in structural and protective applications, yet their viscoelastic response across experimentally relevant time and frequency scales remains difficult to describe consistently Colmenero (2015); Wang et al. (2024). High-frequency molecular simulations (GHz–THz) provide insight into relaxation dynamics near the glass transition, but low-frequency behavior (Hz) relevant for load-bearing applications ASTM (2013); Lakes (2009, 2017) are typically modeled through phenomenological frameworks Tschoegl (2012) that compress broad relaxation spectra into effective parameters. As a result, establishing a direct molecular-to-macroscopic link in polymer-glass mechanics remains a central challenge.
Experimentally, distinct techniques probe disjoint frequency windows: DMA accesses mHz–Hz Menard and Menard (2020); Müller-Pabel et al. (2022); Henriques et al. (2018); Zaccone (2023); Wyss et al. (2007); Sun et al. (2025); Mulliken and Boyce (2006); Split-Hopkinson tests reach the kHz regime Mulliken and Boyce (2006); Sahraoui and Lataillade (1994); Moy et al. (2003); ultrasonic spectroscopy spans MHz Afifi (2003); Afifi and Hasan (2003); and Brillouin scattering extends into the GHz regime Merklein et al. (2022); Weishaupt et al. (1995); Orr et al. (2025), where recent work shows consistency with atomistic modeling Delasoudas et al. (2025). Although these methods reveal different segments of the relaxation spectrum, they do not independently provide a unified microscopic picture.
On the computational side, elastic moduli of disordered solids have been explored through stress-fluctuation approaches and related formalisms Barrat et al. (1988); Leonforte et al. (2005); Procaccia et al. (2016); Kriuchevskyi et al. (2017b); Krief and Ashkenazy (2021); Theodorou and Suter (1986); Lempesis et al. (2013); Vogiatzis et al. (2020); Wittmer et al. (2013, 2015); Kriuchevskyi et al. (2017a), activation–relaxation theories Dyre et al. (1996); Olsen et al. (1998), parallel-replica dynamics Perez et al. (2015), and non-affine approaches Léonforte et al. (2006); Zaccone (2023); Zaccone and Scossa-Romano (2011); Vaibhav et al. (2024); Singh et al. (2025). However, direct MD remains limited to a narrow, high-frequency window due to computational constraints Mukherji et al. (2025).
Non-affine lattice dynamics (NALD) offers a microscopic route to connect atomic structure and macroscopic elasticity Elder et al. (2019); Vaibhav et al. (2024). In NALD, structural disorder generates unbalanced affine forces that drive non-affine relaxations Zaccone (2023), producing a negative correction to the Born modulus Zaccone and Scossa-Romano (2011). The vibrational density of states and affine force correlator together determine the frequency-dependent modulus. However, previous formulations typically assumed a constant (Markovian) memory kernel, whereas glassy polymers exhibit broad, non-Markovian relaxation spectra requiring memory-dependent friction Milster et al. (2024); Li et al. (2017).
Here we extend NALD by introducing a power-law memory kernel within the Generalized Langevin Equation, enabling the description of secondary relaxations. We apply this framework to PMMA, a well-studied polymer glass Mulliken and Boyce (2006); Richeton et al. (2007); Ionita et al. (2015); Hu et al. (2016); Lu et al. (1997); Bodé et al. (2022), using atomistic simulations to compute the vibrational modes Wales and Bogdan (2006); Keyes (1994); Bembenek and Laird (1995); Stratt (1995); Zhang et al. (2019); Vaibhav et al. (2025) and affine force correlator. The resulting shear modulus spans more than twenty decades in frequency, from terahertz vibrational dynamics to millihertz mechanical response, and agrees with oscillatory-shear MD, Brillouin spectroscopy Weishaupt et al. (1995), ultrasonic data Afifi (2003), Split-Hopkinson tests Mulliken and Boyce (2006); Sahraoui and Lataillade (1994); Moy et al. (2003), and our DMA measurements, consistent with the known presence of local relaxations Ngai (2011); Mulliken and Boyce (2006); Moy et al. (2003); Sahraoui and Lataillade (1994). Supported by analytic low-frequency scaling of $\Gamma(\omega)$ and $g(\omega)$ Zaccone and Scossa-Romano (2011); Milkus and Zaccone (2017); Palyulin et al. (2018), the method bridges optical and mechanical regimes and reveals the full multiscale relaxation spectrum of PMMA, with implications for glass-transition modeling Glova and Karttunen (2024). A key result of this work is a description of polymer viscoelasticity across timescales of conventional mechanical testing (DMA), high rate tests (Split-Hopkinson testing, ultrasonics, and Brillouin spectroscopy), and molecular dynamics, ultimately spanning twenty frequency decades across molecular processes to application-relevant timescales.
@@
For the numerical study, we have considered 64 polymeric chains (each containing 10 monomers) of PMMA in a periodic simulation box, making a system of $N=9920$ atoms. Further details are provided in Section S2 of the Supplementary Information (SI) 1. The atomic interactions are modeled using the General AMBER Force Field (GAFF), where bond and angle terms are harmonic and dihedral interactions are described by periodic potentials. Electrostatic interactions are treated using the particle-particle particle-mesh (PPPM) method, with a real-space cut-off of $9$ Å. Lennard-Jones interactions are truncated at $9$ Å. Large-scale molecular dynamics simulations are carried out using LAMMPS Thompson et al. (2022) to generate polymer configurations under fixed temperature and pressure conditions. The system temperature and pressure are controlled through the Nosé–Hoover thermostat and barostat, respectively, with the pressure maintained at zero throughout the simulation. The equations of motion are solved under periodic boundary conditions using the velocity-Verlet integrator with a timestep of 1 fs.
@@
In NALD, the potential energy surface (PES) Wales and Bogdan (2006) is required to measure the mechanical response of atomistic systems. PES describes the total potential energy of the system as a function of the atomic positions, effectively defining the energy landscape in which the atoms reside. The first derivative of the PES with respect to atomic positions gives the forces acting on each atom, which govern the atomic motion and local stability of the system, including internal stresses. The second derivative, known as the Hessian or dynamical matrix, provides information about the curvature of the PES, which determines the vibrational modes and eigenfrequencies of the system. These vibrational properties are crucial for calculating the affine (infinite-frequency) elastic modulus as well as the non-affine corrections arising from local, disorder-induced atomic rearrangements. We diagonalize the Hessian matrix to obtain the eigenfrequencies and corresponding eigenmodes of the system that are used to calculate the normalized vibrational density of states $g(\omega)=\frac{1}{3N-3}\sum_{i}\delta(\omega-\omega_{i})$ and the affine force field correlator $\Gamma_{\alpha\beta\kappa\chi}(\omega)=\langle\hat{\Xi}_{\alpha\beta}\hat{\Xi}_{\kappa\chi}\rangle_{\omega\in[\omega_{k},\omega_{k}+\delta\omega_{k}]}$ (see section S1 of SI 1 for definition), where $\hat{\Xi}_{\alpha\beta}(\omega)$ is the component of affine force field. The frequency-dependent elastic constants $C_{\alpha\beta\kappa\chi}(\Omega)$ can be obtained in terms of these quantities via the relation Zaccone (2023); Zaccone and Scossa-Romano (2011)
@@
where $\Omega$ is the applied strain frequency, $C^{\mathrm{A}}_{\alpha\beta\kappa\chi}$ is the affine contribution to elasticity, $\rho$ is number density, and $\nu$ is the friction memory kernel. The integrand term accounts for non-affine effects that act as a negative correction to the quasi-static limit, arising from local asymmetries in the atomic potential energy landscape during macroscopically affine deformations, as typically observed in glasses as well as in noncentrosymmetric crystals Cui et al. (2019).
The various components of the elastic constants determined using Eq. (1) depend on different components of the affine force field with the corresponding component of the affine elastic constant. Here, we focus on the shear modulus $G$, for which Eq. (1) can be rewritten in the form
@@
Here, $G^{*}(\Omega)$ represents the complex shear modulus and $G_{A}$ is the affine modulus. We separate the real and imaginary components of $G^{*}(\Omega)$ (SI 1, section S1) that correspond to the storage and loss moduli, respectively, which quantify the elastic and viscous responses of the system. We also absorb the mass term in the affine force correlator and friction memory kernel Kriuchevskyi et al. (2020); Vaibhav et al. (2024) to get corresponding mass-scaled quantity.
For simplicity, a fixed value of memory kernel $\nu$} is generally used in Eq. (2), which is the only parameter adjusted to perfectly match the MD results for the infinite-frequency plateau, as done also in Vaibhav et al. (2024); Elder et al. (2019) and discussed in Section S1 of SI 1. In a Markovian process, friction is treated as a constant, as is typically assumed in molecular simulations using, for example, the Langevin thermostat. However, in real glassy materials, friction is more accurately described within a non-Markovian framework, where the memory kernel depends on time or the history of deformation. It is possible to calculate the memory kernel using either the velocity autocorrelation function (VACF) or the force autocorrelation function through a reconstruction technique Koch et al. (2024); Jung et al. (2017); Li et al. (2017). In polymeric systems, the memory kernel typically exhibits a power-law behavior at intermediate and long times Mahato et al. (2025). To capture experimental conditions, we allow the friction parameter $\nu$ to vary and adopt a simple power-law decay Milster et al. (2024); Li et al. (2017) in the time domain, ${\nu}(t)\sim\nu_{0}\,t^{\delta-1}$ with $\delta<1$. In this work we use the memory kernel in the frequency domain as the corresponding Fourier transform that gives ${\nu}(\Omega)={\nu}_{0}\Omega^{-\delta}$, indicating that the effective friction decreases at high frequencies and recovers the constant Markovian limit as $\delta\to 0$. We use this memory kernel in the NALD equations to calculate the shear modulus in order to compare with the experimental results.
@@
We employ the NALD method to calculate the viscoelastic response and compare the results with experimental measurements. For the NALD calculations, we require microscopic information, namely the vibrational density of states $g(\omega)$, the affine force field correlator $\Gamma(\omega)$, and the friction memory kernel $\nu$. In Fig. 1(a) and (b), we plot the vibrational density of states (vDOS) and the affine force field correlator $\Gamma(\omega)$, respectively, as functions of the normal mode frequency $(\omega)$ at room temperature, T=300T=300K. Since our system is at a finite temperature, the vDOS contains a significant fraction of imaginary modes (instantaneous normal modes, INMs) (conventionally represented on the negative branch of $\omega$) in addition to real modes Keyes (1994); Bembenek and Laird (1995); Stratt (1995); Zhang et al. (2019); Vaibhav et al. (2025). The imaginary INMs originate from the presence of saddle points in the potential energy landscape, whereas the real modes correspond to stable vibrational motions (e.g. phonons). The vDOS calculated from the Fourier transform of VACF Mukherji (2024) and as well as from neutron scattering experiments Szymoniak et al. (2025), only provide the behaviour in the real-frequency domain.
@@
We observe several peaks across the entire frequency spectrum. These features may arise from the characteristic bonding nature of PMMA, where certain bonding interactions contribute vibrational modes in the high-frequency range of the vDOS. Near zero frequency, on both the positive and negative branches, the vDOS increases linearly, as often observed in glasses and supercooled liquids Phillips et al. (1989); Bembenek and Laird (1995); Palyulin et al. (2018). This low-frequency regime shows only little variation with temperature (as shown in Fig. S1 of SI 1).
@@
In Fig. 1(b), the affine-force field correlator $\Gamma(\omega)$ is shown for the same temperature. The overall profile of $\Gamma(\omega)$ exhibits several peaks, similar to the vDOS, with peak intensities increasing modestly as the temperature decreases (SI 1, section S1). At high frequencies, around $550\,\mathrm{THz}$, a pronounced peak is observed; however, its influence on the shear modulus is minimal. In contrast, at low frequencies, $\Gamma(\omega)$ follows the expected analytical scaling  Zaccone and Scossa-Romano (2011); Milkus and Zaccone (2017) $\Gamma(\omega)\sim\omega^{2}$, which plays a dominant role in the determination of the modulus. This low-frequency behavior reflects the long-wavelength vibrational modes that couple strongly with the macroscopic deformation response.
@@
Using the information from the vibrational density of states $g(\omega)$ and the affine force field correlator $\Gamma(\omega)$, we calculate the shear modulus as a function of the external frequency $\Omega$ at different temperatures below the glass transition. An important additional quantity in this calculation is the friction memory kernel, which enters the equations in the power-law form Milster et al. (2024); Li et al. (2017) $\nu(\Omega)=\nu_{0}\Omega^{-\delta}$ with parameters $\nu_{0}=5.2\times 10^{18}\,\mathrm{kg\,s^{-1}}$ and $\delta=0.35$. In the time domain, this form corresponds to the memory kernel given by $\nu(t)\sim\nu_{0}t^{(\delta-1)}$. We plot and discuss $\nu$ in the SI 1 as Figure S4 shows $\nu(\Omega)$ as a function of the external frequency.
@@
In the determination of the shear modulus $G$, we perform the integration over both the real and imaginary branches of $\omega$, as shown in Eqs. (S.4, S.5) of SI 1. Since the system is relatively small, finite-size effects arise, leading to the presence of non-physical low-frequency modes. To address this, a fraction of the small-frequency modes below a cut-off frequency $\omega_{\rm min}$ is discarded during the integration, as the system cannot support propagating modes with frequencies lower than the speed of sound in the medium. The minimum frequency can be estimated as $\omega_{\rm min}=\frac{2\pi}{L}\sqrt{\frac{G}{\rho}}$, where $L$ is the size of the simulation box, $\rho$ is the mass density, and $G$ is the shear modulus Vaibhav et al. (2024). Hence, when performing the finite analytical integration for the modulus calculations, we incorporate all modes with $|\omega|<\omega_{\min}$ using analytical theory, to account for the low-frequency contributions below the cut-off (cf. Fig. S2 in the SI 1).
@@
We observe a secondary $\beta$ relaxation Ngai (2011) in the shear modulus in the frequency domain around 1 Hz. This is reflected in the time domain, where $G(t)$ decays on a time-scale $\tau\sim 1$ s. Capturing such small frequencies in simulations is very challenging due to system size limitations and the computational cost of Hessian diagonalization. To overcome this, we use an analytical-theory extension in the very low-frequency regime, exploiting the known behavior of $\Gamma(\omega)$ and $g(\omega)$ Zaccone (2023). Specifically, we fit $\Gamma(\omega)=a_{1}\,\omega^{2},\quad g(\omega)=a_{0}\,\omega,$ with $a_{1}=3.1\times 10^{-18}$ and $a_{0}=4.5\times 10^{-16}$ as shown in Figs. S2 and S3 of the SI 1. We then evaluate the additional analytical contribution, obtaining the exact integral solution at finite integration limits for both the real and imaginary parts of the eigenfrequencies. In this regime, the friction kernel is found to be on the order of $\nu=5\times 10^{22}\,\mathrm{kg\,s^{-1}}$ across the entire deformation frequency range.
@@
Figure 2: Temperature dependence of shear storage modulus $G^{\prime}$ for PMMA calculated using NALD (closed triangles) is compared with DMA experimental data (closed diamonds, at external frequency $\Omega=1$ Hz). The solid line represents the theoretical fit based on the equation described in SI 1, with a glass transition temperature of $T_{g}\approx 390~\mathrm{K}$. We also plot the affine modulus $G_{A}$ (open squares) as a function of temperature.
@@
In Fig. 2, we show the shear modulus $G^{\prime}$ serving as an order parameter that distinguishes the glassy and liquid phases. Here, $G^{\prime}$ is plotted as a function of temperature at a very low deformation frequency of 1 Hz. We compare the experimental DMA results (closed diamonds) with atomistic NALD calculations (closed triangles), which show a reasonable agreement. We also fit the experimental data using the theoretical Zaccone-Terentjev expression for $G^{\prime}$ vs $T$ (details in SI 1) Zaccone and Terentjev (2013), capturing the sharp decay near the glass transition temperature $T_{g}=390~\mathrm{K}$ Ruscher et al. (2019); Mohammadi et al. (2017); Chimenti et al. (2023). Although the NALD calculations capture the slow decrease in the modulus below $T_{g}$, the sharp decay near $T_{g}$ is not well captured. Moreover, we observed a slight decrease in the affine modulus $G_{A}$ near and above $T_{g}$ due to enhanced monomer mobility and reduced topological constraints.
@@
Figure 3: Low-frequency shear storage modulus $G^{\prime}$ as a function of external deformation frequency $\Omega$ at temperature $T=300$ K. The Solid line represents NALD calculations and symbols represent MD and experimental results that include DMA study in this work and other experiments from previous studies Weishaupt et al. (1995); Fukuhara and Sampei (1995); Afifi (2003); Afifi and Hasan (2003); Mott et al. (2008); Sahraoui and Lataillade (1994); Davies and Hunter (1963); Moy et al. (2003); Mulliken and Boyce (2006); Christöfl et al. (2021).